OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL A class of 21 students has 9 girls and 12 boys. What is the probability of a teacher randomly calling on a boy

OpenStudy (anonymous):

wait for the answers plzz

OpenStudy (anonymous):

57 i think

OpenStudy (anonymous):

\[A. \frac{ 3 }{ 4}\] \[B. \frac{ 4 }{ 7 }\] \[C.\frac{ 3 }{7}\] \[D. \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

@xGuardians

OpenStudy (anonymous):

B I think

OpenStudy (xguardians):

12 (# of boys)/21(# of total students) 12/21 = 4/7 (Simplified) B.